# Multi-Crit: Benchmarking Multimodal Judges on Pluralistic Criteria-Following

### v1

@@ 3.5 Measuring Pluralistic Criteria Following
Pluralistic Accuracy (PAcc).
Measures whether the judge produces correct judgments across all criteria for each evaluation instance. Let $X$ denote the set of evaluation instances, where each $x=(q,l_{a},l_{b})\in X$ represents a multimodal prompt with its paired candidate responses, and let $C_{x}$ denote the set of evaluation criteria associated with $x$:
@@
Trade-off Sensitivity (TOS).
Measures whether the judge model recognizes criterion-level trade-offs within each evaluation instance. Let $\mathcal{P}_{x}=\{\,(c_{i},c_{j})\in C_{x}\times C_{x}\mid y_{x,c_{i}}\neq y_{x,c_{j}}\,\}$ denotes the set of criterion pairs in instance $x$ that exhibit ground-truth conflicts.
@@
Conflict Matching Rate (CMR).
Measures the judge’s ability to correctly resolve ground-truth conflicts between criterion-based judgments. A conflict is considered matched only if the model predicts both criteria in agreement with the ground truth preference:
@@ 4 Experiments
To comprehensively assess multimodal judge boundaries, we evaluate 25 LMMs across three model groups: (1) proprietary models, including the GPT series [hurst2024gpt, jaech2024openai, openai2025o3o4mini], Gemini family [comanici2025gemini], and Claude-3.7-Sonnet [anthropic_claude3.7_sonnet_2025]; (2) leading open-source LMMs, including Qwen2.5/3-VL [bai2025qwen2, yang2025qwen3], InternVL3/3.5 [zhu2025internvl3, wang2025internvl3.5], GLM-4.1V-Thinking [hong2025glm], LLaMA3.2-V [grattafiori2024llama], MiniCPM-V-4.5 [yu2025minicpm], Eagle2.5 [chen2025eagle], Molmo [deitke2025molmo] and LLaVA-OneVision [li2024llava]; and (3) finetuned judge models, including LLaVA-Critic [xiong2025llava]/-R1 [wang2025llava], R1-Reward [zhang2025r1] and UnifiedReward [wang2025unified].
Implementation Details.
@@
During inference, each model was prompted to evaluate a single multimodal sample containing one prompt and a pair of responses, focusing exclusively on one criterion at a time. The full evaluation prompts are provided in the Appendix, with minor adjustments for model-specific formatting. Sampling temperatures follow the original repository settings, or default to 0.6 with a top-$p$ of 0.95 when unspecified.
@@
Model Prompt-Level Criterion-Level Pluralisic Conflict Tradeoff Grounding Logic Hallucination Exploration Efficiency Avg Open-Source LMMs LLaVA-OneVision-7B 9.52 17.08 51.38 40.74 44.19 29.31 50.55 58.72 44.70 Llama-3.2-11B-Vision-Instruct 10.32 19.93 44.95 49.38 45.35 43.10 53.85 62.39 50.81 Molmo-7B 11.11 13.17 34.86 45.68 40.70 43.10 43.96 61.47 46.98 Qwen3-VL-8B-Instruct 16.67 19.22 35.78 50.62 56.98 65.52 47.25 64.22 56.92 Qwen2.5-VL-7B-Instruct 16.67 25.62 66.06 46.91 52.33 55.17 57.14 65.14 55.34 Eagle2.5-8B 19.05 27.05 48.62 49.38 47.67 44.83 58.24 69.72 53.97 GLM-4.1V-9B-Thinking 23.02 27.40 55.05 54.32 60.47 65.52 50.55 66.97 59.57 MiniCPM-V-4.5-8B 25.40 38.08 75.23 56.79 66.28 56.90 65.93 66.97 62.57 InternVL3-8B-Instruct 26.98 39.50 66.06 64.20 65.12 58.62 67.03 76.15 66.22 InternVL3-78B-Instruct 29.37 39.50 67.89 62.96 69.77 53.45 63.74 74.31 64.85 Qwen2.5-VL-72B-Instruct 32.54 45.91 77.06 60.49 61.63 56.90 68.13 75.23 64.48 InternVL3.5-8B 32.54 39.15 69.72 65.43 63.95 56.90 71.43 71.56 65.85 MiMo-VL-7B 37.30 41.99 71.56 59.26 59.30 70.69 67.03 75.23 66.30 InternVL3.5-38B 37.30 47.69 75.23 71.60 70.93 55.17 78.02 73.39 69.82 Finetuned Judges UnifiedReward-7B (Qwen2.5-VL) 11.90 13.17 22.02 64.2.0 70.93 50.00 58.24 54.13 59.50 LLaVA-Critic-7B (LLaVA-OV) 13.49 22.42 50.46 43.21 44.19 43.10 45.05 60.55 47.22 LLaVA-Critic-R1-7B(Qwen2.5-VL) 16.67 16.37 49.54 53.09 50.00 51.72 56.04 55.05 53.18 R1-Reward (Qwen2.5-VL) 19.05 24.56 62.39 58.02 54.65 46.55 52.75 60.55 54.50 Proprietary Judge Models Claude-3.7-Sonnet 36.51 51.96 83.49 74.07 62.79 65.52 68.13 76.15 69.33 Gemini-2.5-Flash 37.30 53.05 73.39 67.50 63.95 64.91 78.02 77.98 70.47 Gemini-2.5-Pro 41.27 52.33 75.93 79.01 67.44 68.97 75.56 74.31 73.06 GPT-4o 41.27 55.16 84.40 61.73 67.44 55.17 80.22 84.40 69.79 o3 44.44 62.28 82.57 67.90 70.93 84.21 84.62 81.65 77.86 GPT-5 45.24 56.58 78.90 74.07 70.93 82.76 81.32 77.98 77.41 o4-mini 53.17 65.84 83.49 77.78 75.58 79.31 83.52 88.07 80.85
@@ 4.1 Main Results
Pluralistic judging is highly challenging, especially for open-ended tasks.
The Pluralistic Accuracy (PAcc), which requires correct preferences across all criteria for a single query, is notably low. o4-mini and Claude-3.7-Sonnet achieve the highest open-ended PAcc at 32.78% and 31.77%, while o4-mini and GPT-5 lead on verifiable reasoning with 53.17% and 45.24%. This highlights the difficulty of pluralistic judging even for state-of-the-art LMMs. As in Figure 3, for top-performing models, results on open-ended tasks are lower, reflecting the greater subjectivity and reliance on fine-grained visual perception for evaluating free-form responses (e.g., real-world images or screenshots). In contrast, existing multimodal reasoning tasks—typically involving multimodal math, STEM, puzzle, and coding problems—are more objective, with clearer verification and explicit visual grounding. These findings highlight the inherent difficulty of pluralistic judgment and underscore the necessity of the proposed Multi-Crit benchmark.
Model performance varies across criteria and metrics.
For judgment in verifiable reasoning, o4-mini performs best on Logic (75.58%) and Efficiency (88.07%), but is surpassed by o3 in Hallucination (84.21%) and Exploration (84.62%), and by Gemini-2.5-Pro in Grounding (79.01%). At the prompt level, GPT-4o shows the highest sensitivity to trade-offs in prompts with criterion conflicts, achieving the best Trade-off Sensitivity of 84.40%. However, its ability to resolve conflicting criterion-level preference pairs is notably weaker than that of the top-performing o4-mini (55.16% vs. 65.84% in CMR). A similar trend is observed in open-ended tasks: GPT-4o- excels in Expressiveness (76.17%) but lags in Completeness (68.14%) and Hallucination (65.75%). In contrast, o3 performs strongly on Hallucination (72.93%) but ranks low in Expressiveness (63.21%). These results highlight the multifaceted nature of pluralistic judgment—no model excels across all aspects, each showing distinct strengths and blind spots.
Open-source LMMs fall short in pluralistic criterion-following, especially in capturing criterion-level preference conflicts.
As shown in Tables 3 and 4, proprietary LMMs (e.g., o4-mini, GPT-5, o3) maintain a clear lead over open-source counterparts (e.g., InternVL3.5-38B, Qwen2.5-VL-72B-Instruct) across all prompt-level metrics (PAcc, CMR, and TOS) and criterion-level accuracies. Specifically, the conflict matching rate drops by about 9.4 points in open-ended tasks (43.1 $\rightarrow$ 33.7) and 18.1 points in reasoning tasks (65.8 $\rightarrow$ 47.7), far exceeding the 4- and 11-point declines observed in criterion-level accuracy. This indicates that open-source models struggle to maintain consistent pluralistic evaluation across conflicting criteria, highlighting their limited ability to integrate multi-faceted criterion-level signals into reliable judgments.
Finetuned judge models improve evaluation in visual grounding but are limited in handling pluralistic criteria.
@@ 4.2 Additional Discussions
1. Does RLVR on multimodal reasoning improve judgment in reasoning traces?
We evaluate three GRPO-finetuned models [wang2025sota, meng2025mm, chen2025sftrlearlyinvestigation] on general reasoning tasks (e.g., math and charts), all showing clear reasoning gains. However, these models, trained on holistic accuracy rewards, show reduced sensitivity to trade-offs and conflict matching in reasoning judgments (Figure 4), indicating weaker recognition of criterion-level preference conflicts.
@@ 2. Does test-time scaling hold?
3. What are the current boundaries of criteria following for open-source and proprietary models?
To probe the upper limits of criterion-level consistency, we correlate each model group’s strongest criterion accuracies with human inter-annotator agreement (Cohen’s $\kappa$), excluding reasoning-efficiency, which is largely influenced by response length. As in Figure 6, proprietary models exhibit a strong and significant correlation (r=0.73, p=0.024), indicating that their upper-bound judgment patterns align closely with human agreement. In contrast, open-source models show a weaker, more inconsistent relationship (r=0.36, p=0.344), reflecting limited capacity to internalize human-consistent evaluation criteria. These findings suggest that while open-source models would benefit from scaling high-quality human annotations to improve pluralistic criteria-following, proprietary systems might face the next challenge of surpassing human-level evaluation alignment.
@@ C.1 Additional Benchmark Details
Models for Generating Candidate Responses:
Gemini-2.5-Pro [comanici2025gemini], Gemini-2.5-Flash [comanici2025gemini], GPT-4.1 [openai2024gpt4_1], GPT-4o [achiam2023gpt], Qwen2.5-VL-7B-Instruct [bai2025qwen2], Qwen2.5-VL-72B-Instruct [bai2025qwen2], LLaMA-3.2-11B-Vision-Instruct [grattafiori2024llama], InternVL3-8B-Instruct [zhu2025internvl3], InternVL3-38B-Instruct [zhu2025internvl3], MiMo-VL-7B-RL [team2025kimi], GLM-4.1V-9B-Thinking [hong2025glm].
Specifically, for prompts from WildVision-Battle [lu2024wildvision], if the two original responses receive a tie in their human preference, we keep the original pair; otherwise, we generate new responses using the models listed above.
Detailed Evaluation Criteria.
Table 7 presents the detailed criteria for judging open-ended generation, and Table 8 lists those for judging verifiable reasoning. To better guide both human annotators and judge LMMs, we organize the description of each criterion into three components:
@@
Evaluation Prompt Template.
We provide the evaluation prompt used for judge-model inference in Table 9. When evaluating LMM judges on the Multi-Crit benchmark, each criterion instance—consisting of the question, the pair of responses, and the target criterion—is filled into this template and sent to the model during inference. For finetuned judges [xiong2025llava, wang2025llava, zhang2025r1, wang2025unified], we adapt this template to match the prompt format used in their original repositories, avoiding significant out-of-distribution judgment behaviors.
Benchmark Resp. Quantile Multi-Criterion Anno.
MLLM-as-a-Judge [chen2024mllm] (54, 89, 153)
VL-RewardBench [li2025vl] (48, 99, 136)
MM-RLHF [zhang2025mm] (16, 49, 144)
Multimodal-RewardBench [yasunaga2025multimodal] (1, 72, 138)
Multi-Crit (ours) (104, 164, 247)
Table 5: Comparison with existing benchmarks.
@@
Table 7: Five evaluation criteria for judging open-ended content generation tasks in Multi-Crit Benchmark.
@@
C.2 Additional Data Statistics
Response Pair Composition.
As shown in Table 6, for both splits, roughly $40\%$ of response pairs originate from the same model via random sampling, while the remaining pairs are generated by different models. This balanced mix ensures that Multi-Crit evaluates whether LMM judges can capture response-quality differences both across models (reflecting model and training diversity) and within the same model (reflecting sampling diversity).
Cross-Criterion Correlation and Conflict Frequency.
We report additional statistics on criterion-level human preferences to illustrate how Multi-Crit captures cross-criterion correlations and conflicts. Figure 7 shows how human preferences correlate across criteria on prompts that exhibit criterion-level conflicts. Figure 8 reports how often two criteria yield opposite human preferences on the same response pair. Multi-Crit contains preference conflicts across all criterion pairs in both splits, reflecting both the nuanced differences within the response pairs and the inherent diversity and tension across evaluation dimensions. Certain criterion pairs display intuitive trade-offs—such as completeness vs. no-hallucination and no-hallucination vs. expressiveness in the open-ended split, and exploration vs. efficiency in the reasoning split—and these pairs also exhibit the highest frequencies of human-annotated conflicts.

### v2

@@ 3.5 Measuring Pluralistic Criteria Following
Pluralistic Accuracy (PAcc). Measures whether the judge produces correct judgments across all criteria for each evaluation instance. Let $X$ denote the set of evaluation instances, where each $x=(q,l_{a},l_{b})\in X$ represents a multimodal prompt with its paired candidate responses, and let $C_{x}$ denote the set of evaluation criteria associated with $x$:
@@
Trade-off Sensitivity (TOS). Measures whether the judge model recognizes criterion-level trade-offs within each evaluation instance. Let $\mathcal{P}_{x}=\{\,(c_{i},c_{j})\in C_{x}\times C_{x}\mid y_{x,c_{i}}\neq y_{x,c_{j}}\,\}$ denotes the set of criterion pairs in instance $x$ that exhibit ground-truth conflicts.
@@
Conflict Matching Rate (CMR). Measures the judge’s ability to correctly resolve ground-truth conflicts between criterion-based judgments. A conflict is considered matched only if the model predicts both criteria in agreement with the ground truth preference:
@@ 4 Experiments
To comprehensively assess multimodal judge boundaries, we evaluate 25 LMMs across three model groups: (1) proprietary models, including the GPT series [16, 17, 31], Gemini family [8], and Claude-3.7-Sonnet [32]; (2) leading open-source LMMs, including Qwen2.5/3-VL [2, 55], InternVL3/3.5 [73, 43], GLM-4.1V-Thinking [14], LLaMA3.2-V [11], MiniCPM-V-4.5 [59], Eagle2.5 [4], Molmo [9] and LLaVA-OneVision [21]; and (3) finetuned judge models, including LLaVA-Critic [53]/-R1 [45], R1-Reward [66] and UnifiedReward [48].
@@
Implementation Details. During inference, each model was prompted to evaluate a single multimodal sample containing one prompt and a pair of responses, focusing exclusively on one criterion at a time. The full evaluation prompts are provided in the Appendix, with minor adjustments for model-specific formatting. Sampling temperatures follow the original repository settings, or default to 0.6 with a top-$p$ of 0.95 when unspecified.
@@
Model Prompt-Level Criterion-Level Pluralisic Conflict Tradeoff Grounding Logic Hallucination Exploration Efficiency Avg Open-Source LMMs LLaVA-OneVision-7B 9.52 17.08 51.38 40.74 44.19 29.31 50.55 58.72 44.70 Llama-3.2-11B-Vision-Instruct 10.32 19.93 44.95 49.38 45.35 43.10 53.85 62.39 50.81 Molmo-7B 11.11 13.17 34.86 45.68 40.70 43.10 43.96 61.47 46.98 Qwen3-VL-8B-Instruct 16.67 19.22 35.78 50.62 56.98 65.52 47.25 64.22 56.92 Qwen2.5-VL-7B-Instruct 16.67 25.62 66.06 46.91 52.33 55.17 57.14 65.14 55.34 Eagle2.5-8B 19.05 27.05 48.62 49.38 47.67 44.83 58.24 69.72 53.97 GLM-4.1V-9B-Thinking 23.02 27.40 55.05 54.32 60.47 65.52 50.55 66.97 59.57 MiniCPM-V-4.5-8B 25.40 38.08 75.23 56.79 66.28 56.90 65.93 66.97 62.57 InternVL3-8B-Instruct 26.98 39.50 66.06 64.20 65.12 58.62 67.03 76.15 66.22 InternVL3-78B-Instruct 29.37 39.50 67.89 62.96 69.77 53.45 63.74 74.31 64.85 Qwen2.5-VL-72B-Instruct 32.54 45.91 77.06 60.49 61.63 56.90 68.13 75.23 64.48 InternVL3.5-8B 32.54 39.15 69.72 65.43 63.95 56.90 71.43 71.56 65.85 MiMo-VL-7B 37.30 41.99 71.56 59.26 59.30 70.69 67.03 75.23 66.30 InternVL3.5-38B 37.30 47.69 75.23 71.60 70.93 55.17 78.02 73.39 69.82 Finetuned Judges UnifiedReward-7B (Qwen2.5-VL) 11.90 13.17 22.02 64.20 70.93 50.00 58.24 54.13 59.50 LLaVA-Critic-7B (LLaVA-OV) 13.49 22.42 50.46 43.21 44.19 43.10 45.05 60.55 47.22 LLaVA-Critic-R1-7B(Qwen2.5-VL) 16.67 16.37 49.54 53.09 50.00 51.72 56.04 55.05 53.18 R1-Reward (Qwen2.5-VL) 19.05 24.56 62.39 58.02 54.65 46.55 52.75 60.55 54.50 Proprietary Judge Models Claude-3.7-Sonnet 36.51 51.96 83.49 74.07 62.79 65.52 68.13 76.15 69.33 Gemini-2.5-Flash 37.30 53.05 73.39 67.50 63.95 64.91 78.02 77.98 70.47 Gemini-2.5-Pro 41.27 52.33 75.93 79.01 67.44 68.97 75.56 74.31 73.06 GPT-4o 41.27 55.16 84.40 61.73 67.44 55.17 80.22 84.40 69.79 o3 44.44 62.28 82.57 67.90 70.93 84.21 84.62 81.65 77.86 GPT-5 45.24 56.58 78.90 74.07 70.93 82.76 81.32 77.98 77.41 o4-mini 53.17 65.84 83.49 77.78 75.58 79.31 83.52 88.07 80.85
@@ 4.1 Main Results
Pluralistic judging is highly challenging, especially for open-ended tasks. The Pluralistic Accuracy (PAcc), which requires correct preferences across all criteria for a single query, is notably low. o4-mini and Claude-3.7-Sonnet achieve the highest open-ended PAcc at 32.78% and 31.77%, while o4-mini and GPT-5 lead on verifiable reasoning with 53.17% and 45.24%. This highlights the difficulty of pluralistic judging even for state-of-the-art LMMs. As in Figure 3, for top-performing models, results on open-ended tasks are lower, reflecting the greater subjectivity and reliance on fine-grained visual perception for evaluating free-form responses (e.g., real-world images or screenshots). In contrast, existing multimodal reasoning tasks—typically involving multimodal math, STEM, puzzle, and coding problems—are more objective, with clearer verification and explicit visual grounding. These findings highlight the inherent difficulty of pluralistic judgment and underscore the necessity of the proposed Multi-Crit benchmark.
Model performance varies across criteria and metrics. For judgment in verifiable reasoning, o4-mini performs best on Logic (75.58%) and Efficiency (88.07%), but is surpassed by o3 in Hallucination (84.21%) and Exploration (84.62%), and by Gemini-2.5-Pro in Grounding (79.01%). At the prompt level, GPT-4o shows the highest sensitivity to trade-offs in prompts with criterion conflicts, achieving the best Trade-off Sensitivity of 84.40%. However, its ability to resolve conflicting criterion-level preference pairs is notably weaker than that of the top-performing o4-mini (55.16% vs. 65.84% in CMR). A similar trend is observed in open-ended tasks: GPT-4o- excels in Expressiveness (76.17%) but lags in Completeness (68.14%) and Hallucination (65.75%). In contrast, o3 performs strongly on Hallucination (72.93%) but ranks low in Expressiveness (63.21%). These results highlight the multifaceted nature of pluralistic judgment—no model excels across all aspects, each showing distinct strengths and blind spots.
Open-source LMMs fall short in pluralistic criterion-following, especially in capturing criterion-level preference conflicts. As shown in Tables 3 and 4, proprietary LMMs (e.g., o4-mini, GPT-5, o3) maintain a clear lead over open-source counterparts (e.g., InternVL3.5-38B, Qwen2.5-VL-72B-Instruct) across all prompt-level metrics (PAcc, CMR, and TOS) and criterion-level accuracies. Specifically, the conflict matching rate drops by about 9.4 points in open-ended tasks (43.1 $\rightarrow$ 33.7) and 18.1 points in reasoning tasks (65.8 $\rightarrow$ 47.7), far exceeding the 4- and 11-point declines observed in criterion-level accuracy. This indicates that open-source models struggle to maintain consistent pluralistic evaluation across conflicting criteria, highlighting their limited ability to integrate multi-faceted criterion-level signals into reliable judgments.
@@ 4.2 Additional Discussions
1. Does RLVR on multimodal reasoning improve judgment in reasoning traces? We evaluate three GRPO-finetuned models [46, 28, 5] on general reasoning tasks (e.g., math and charts), all showing clear reasoning gains. However, these models, trained on holistic accuracy rewards, show reduced sensitivity to trade-offs and conflict matching in reasoning judgments (Figure 4), indicating weaker recognition of criterion-level preference conflicts.
@@
3. What are the current boundaries of criteria following for open-source and proprietary models? To probe the upper limits of criterion-level consistency, we correlate each model group’s strongest criterion accuracies with human inter-annotator agreement (Cohen’s $\kappa$), excluding reasoning-efficiency, which is largely influenced by response length. As in Figure 6, proprietary models exhibit a strong and significant correlation (r=0.73, p=0.024), indicating that their upper-bound judgment patterns align closely with human agreement. In contrast, open-source models show a weaker, more inconsistent relationship (r=0.36, p=0.344), reflecting limited capacity to internalize human-consistent evaluation criteria. These findings suggest that while open-source models would benefit from scaling high-quality human annotations to improve pluralistic criteria-following, proprietary systems might face the next challenge of surpassing human-level evaluation alignment.
@@ 5 Conclusion
Acknowledgement.
We thank Chunyuan Li for his valuable feedback on the project idea. This work was partially supported by NSF IIS 2347592, 2348169, DBI 2405416, CCF 2348306, CNS 2347617, RISE 2536663.
@@ Appendix B Comparison Against Prior Benchmarks
Benchmark Resp. Quantile Multi-Criterion Anno.
MLLM-as-a-Judge [3] (54, 89, 153)
VL-RewardBench [23] (48, 99, 136)
MM-RLHF [68] (16, 49, 144)
Multimodal-RewardBench [56] (1, 72, 138)
Multi-Crit (ours) (104, 164, 247)
Table 5: Comparison with existing benchmarks.
@@ C.1 Additional Benchmark Details
Models for Generating Candidate Responses: Gemini-2.5-Pro [8], Gemini-2.5-Flash [8], GPT-4.1 [30], GPT-4o [16], Qwen2.5-VL-7B-Instruct [2], Qwen2.5-VL-72B-Instruct [2], LLaMA-3.2-11B-Vision-Instruct [11], InternVL3-8B-Instruct [73], InternVL3-38B-Instruct [73], MiMo-VL-7B-RL [39], GLM-4.1V-9B-Thinking [14].
Specifically, for prompts from WildVision-Battle [27], if the two original responses receive a tie in their human preference, we keep the original pair; otherwise, we generate new responses using the models listed above.
Detailed Evaluation Criteria. Table 7 presents the detailed criteria for judging open-ended generation, and Table 8 lists those for judging verifiable reasoning. To better guide both human annotators and judge LMMs, we organize the description of each criterion into three components:
@@
Evaluation Prompt Template. We provide the evaluation prompt used for judge-model inference in Table 9. When evaluating LMM judges on the Multi-Crit benchmark, each criterion instance—consisting of the question, the pair of responses, and the target criterion—is filled into this template and sent to the model during inference. For finetuned judges [53, 45, 66, 48], we adapt this template to match the prompt format used in their original repositories, avoiding significant out-of-distribution judgment behaviors.
C.2 Additional Data Statistics
Response Pair Composition. As shown in Table 6, for both splits, roughly $40\%$ of response pairs originate from the same model via random sampling, while the remaining pairs are generated by different models. These two types serve complementary roles: cross-model pairs reflect realistic deployment scenarios for automated evaluation (e.g., LMArena or WildVision Arena), where responses differ along multiple dimensions; intra-model pairs control for model identity and surface style, which is crucial when applying judges for test-time scaling and RL reward modeling. Our balanced mix ensures that Multi-Crit assesses whether LMM judges can capture quality differences both across models (reflecting model diversity) and within the same model (reflecting sampling diversity).
@@ C.2 Additional Data Statistics
Cross-Criterion Correlation and Conflict Frequency. We report additional statistics on criterion-level human preferences to illustrate how Multi-Crit captures cross-criterion correlations and conflicts. Figure 7 shows how human preferences correlate across criteria on prompts that exhibit criterion-level conflicts. Figure 8 reports how often two criteria yield opposite human preferences on the same response pair. Multi-Crit contains preference conflicts across all criterion pairs in both splits, reflecting both the nuanced differences within the response pairs and the inherent diversity and tension across evaluation dimensions. Certain criterion pairs display intuitive trade-offs—such as completeness vs. no-hallucination and no-hallucination vs. expressiveness in the open-ended split, and exploration vs. efficiency in the reasoning split—and these pairs also exhibit the highest frequencies of human-annotated conflicts.
@@
Table 7: Five evaluation criteria for judging open-ended generation tasks in Multi-Crit Benchmark.